OpenStudy (mmend98):

The relationship between pressure and moles, when volume and temperature of a gas are held constant, is: P/n = k. We could say then, that: If the number of moles is doubled, the pressure is halved. If the number of moles if halved, the pressure is doubled. If the number of moles is tripled, the pressure also triples. If the number of moles is halved, the pressure is quadrupled.

pooja195 (pooja195):

Sorry. I don't know this so i can't really help. Also this is NOT a math question.

OpenStudy (irishboy123):

\(\huge\frac{P}{n} = const. \implies \\ \\ \huge \frac{k \times P}{k \times n}\ = \ const. \)

OpenStudy (mmend98):

so what answer would u choose here? @irishboy123

OpenStudy (triciaal):